OpenStudy (anonymous):

determine the intercepts: 2x-3y=6

OpenStudy (anonymous):

do i start by using y = mx+b?

OpenStudy (anonymous):

for x-intercept put y=0 and for y-intercept put x=0 and solve the resulting equaions respectively

OpenStudy (anonymous):

? so i graph the line?

OpenStudy (anonymous):

otherwise reduce the given equation to the form (x/a) +(y/b)=1 where a and b are the x and y intercepts respectively

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

2x-3y=6 (2x-3y)=1 2x/6 -3y/6 =1 or x/3 +y/(-2) =1

OpenStudy (anonymous):

ya i got that but from here, how do i graph x/3 +y/2?

OpenStudy (anonymous):

instead of x/3 +y/2 it is x/3 +y/(-2) =1 the line passes through (3,0) and (0,-2)

OpenStudy (anonymous):

but if x/3=0 then it would be x=0*3 and x =0?

OpenStudy (anonymous):

for (x/a) +(y/b)=1 the line passes through (a,0) and (0,b)

OpenStudy (anonymous):

Method 1 for x-intercept we put y=0 in 2x-3y=6 hence we have 2x-3(0)=6 or 2x=6 or x=3

OpenStudy (anonymous):

similarly for y-intercept we put x=0 in 2x-3y=6 we get 2(0)-3y=6 or -3y=6 or y=-2

OpenStudy (anonymous):

double checking hold on

OpenStudy (anonymous):

mmkay i got it from here

OpenStudy (anonymous):

keep it up

OpenStudy (anonymous):

so i got 3,0 and 0,-2

OpenStudy (anonymous):

this the final answer?